The combined SAT scores for the students at a local high school are normally distributed with a mean of 1548 and a standard deviation of 299. The local college includes a minimum score of 2266 in its admission requirements.
What percentage of students from this school earn scores that fail to satisfy the admission requirement?
P(X < 2266) = %
Enter your answer as a percent accurate to 1 decimal place (do not enter the "%" sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
I tried 2266-1548/299

Answer by stanbon(75887)

The combined SAT scores for the students at a local high school are normally distributed with a mean of 1548 and a standard deviation of 299. The local college includes a minimum score of 2266 in its admission requirements.
What percentage of students from this school earn scores that fail to satisfy the admission requirement?
P(X < 2266) = %
Enter your answer as a percent accurate to 1 decimal place (do not enter the "%" sign). Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
z = (2266-1548)/299 = 2.4013
----
P(x < 2266) = P(z < 2.4013) = normalcdf(-100,2.4013) = 0.9918
=================================
Cheers,
Stan H.
==============================